OpenStudy (anonymous):

If a light intensity 67 foot-candles at 15 feet. Find the light intensity at 5 feet

OpenStudy (anonymous):

intensity is inversely proportional to (distance)^2 67/intensity = (5/15)^2

OpenStudy (anonymous):

is 22.333 the right answer?

OpenStudy (anonymous):

No I worked it out and just got 603 which was correct I did 67=k/15^2 k=15075 15075*1/5^2 ANSWER=603